# Hölder spiral arcs

### v2

@@ 1. Introduction
Given a function $\phi:[2\pi,+\infty)\to(0,\infty)$ with $\lim_{t\rightarrow\infty}\phi(t)=0$, we denote by $\mathcal{S}_{\phi}$ the spiral
@@
Let $s>1$ and $\mathcal{S}_{\phi}$ be an almost circular spiral arc. Then $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc if, and only if, $\sum_{n=1}^{\infty}\phi_{n}^{s}$ converges.
@@
Let $0<r\leq p$. There is a $r/p$-Hölder map $f:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$, and a Lipschitz map $g:\mathcal{S}_{r}\rightarrow\mathcal{S}_{p}$. Moreover, every $\alpha$-Hölder map $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ needs to satisfy $\alpha\leq r/p$.
@@
for all $x_{1},x_{2}\in X$. The number $\alpha$ is called the Hölder exponent of $f$, and the smallest $C>0$ is the Hölder semi-norm of $f$, denoted by $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{\alpha}f$.
Recall that a metric space $X$ is a metric arc if there is a homeomorphism $f$ mapping the interval $[0,1]$ onto the space $X$. Given an interval $I\subset[0,1]$, we say that $f(I)$ is a subarc of $X$. Furthermore, if the interval $I$ has endpoints $a,b\in[0,1]$, we say that $f(I\setminus\{a,b\})$ is the interior of the subarc $f(I)$. If $X$ and $f(I)$ have distinct interiors, we say that $f(I)$ is a proper subarc of $X$. If $f$ is $\alpha$-Hölder for some $\alpha\in(0,1)$, we say that $X$ is an $\alpha$-Hölder arc.
@@
2. The $s$-variation and $(1/s)$-Hölder rectifiability
A classification of Lipschitz arcs (and even more general Lipschitz curves) was given by Ważewski in 1927: an arc is a Lipschitz arc if and only if $\mathcal{H}^{1}(X)<\infty$, and if $X$ is a Lipschitz arc, then there exists a $(2\mathcal{H}^{1}(X))$-Lipschitz parameterization $f:[0,1]\to X$. See [AO17, Theorem 4.4] for the proof. An analogous result for $\frac{1}{s}$-Hölder arcs does not exist as there is no connection between $\frac{1}{s}$-Hölder parameterizations and $\mathcal{H}^{s}$; see for example [BNV19, Proposition 9.6]. In this section we give a characterization of $\frac{1}{s}$-Hölder arcs in terms of the $s$-variation, the correct analogue of $\mathcal{H}^{1}$ for $s\geq 1$.
A partition of $X$ is a finite collection of subarcs $\mathcal{P}=\{X_{1},\dots,X_{n}\}$ with disjoint interiors, and with their union being equal to $X$. For $s\geq 1$, define the $s$-variation of $X$ by
@@ 2. The $s$-variation and $(1/s)$-Hölder rectifiability
Letting $\epsilon$ go to 00, we obtain the desired inequality.
@@
The continuity of the function $t\mapsto\|f([0,t])\|_{s\textup{-var}}$ follows by [FV10, Proposition 5.8] applied to the given homeomorphism $f$.
@@
Suppose first that $H_{1/s}(X)<\infty$. Then, there exists a $(1/s)$-Hölder homeomorphism $f:[0,1]\to X$ with $H_{1/s}(X)\leq\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$. Fix arbitrary $H>\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ and $\epsilon\in(0,1)$.
Let $X_{1},\dots,X_{n}$ be a partition of $X$ and for each $i\in\{1,\dots,n\}$ let $I_{i}=f^{-1}(X_{i})$. Then $\{I_{1},\dots,I_{n}\}$ is an interval partition of $[0,1]$. For each $i\in\{1,\dots,n\}$, let $x_{i},y_{i}\in I_{i}$ such that
$|f(x_{i})-f(y_{i})|\geq(1-\epsilon)\operatorname{diam}{f(I_{i})}$
@@
$(\operatorname{diam}{f(I_{i})})^{s}\leq(1-\epsilon)^{-s}|f(x_{i})-f(y_{i})|^{s}\leq(1-\epsilon)^{-s}H^{s}\operatorname{diam}{J_{i}}.$
@@
Taking supremum over all partitions and letting $\epsilon\to 0$ we obtain $\|X\|_{s\text{-var}}\leq H^{s}$. Since $H>\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ is arbitrary, and $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ can be chosen as close to $H_{1/s}(X)$ as necessary, this implies that $\|X\|_{s\text{-var}}\leq H_{1/s}(X)^{s}$.
@@
By Lemma 2.2, $\psi$ is an increasing function from $[0,1]$ into $[0,1]$ and, by Lemma 2.3, $\psi$ is continuous. Since $\psi(0)=0$ and $\psi(1)=1$, it follows that $\psi$ is surjective. By Lemma 2.2(2), it follows that $\psi$ is in fact a homeomorphism. This allows for the definition of $F=f\circ\psi^{-1}:[0,1]\to X$.
It remains to show that $F$ is $(1/s)$-Hölder. Let $0\leq x<y\leq 1$ and let $0\leq x^{\prime}<y^{\prime}\leq 1$ be such that $\psi(x^{\prime})=x$ and $\psi(y^{\prime})=y$. Then,
$|F(x)-F(y)|^{s}$ $=|f(x^{\prime})-f(y^{\prime})|^{s}$
@@
Therefore, $F$ is $(1/s)$-Hölder continuous with $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{F}\leq\|X\|_{s\text{-var}}^{1/s}$. As a result,
$H_{1/s}(X)\leq\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{F}\leq\|X\|_{s\text{-var}}^{1/s}<\infty,$
which allows to apply the first direction and yields $H_{1/s}(X)=\|X\|_{s\text{-var}}^{1/s}$ as needed. This equality and the definition of $H_{1/s}(X)$ complete the proof. ∎
@@ 3. Hölder rectifiability of spiral arcs
Suppose that $\mathcal{S}_{\phi}$ is almost circular. Fix $j\in\mathbb{N}$ and $s\geq 1$. On the one hand, if $h_{j}:[0,1]\to\mathcal{S}_{\phi}^{j}$ is the constant speed Lipschitz map, then for all $x,y\in[0,1]$,
@@
Hence, $H_{1/s}(\mathcal{S}_{\phi}^{j})\simeq\phi_{j}$, and the statement follows directly from Theorem 3.1. ∎
@@
Suppose first that $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc. Fix $k\in\mathbb{N}$, and note that $\mathcal{S}_{\phi}^{1},\dots,\mathcal{S}_{\phi}^{k},\mathcal{S}_{\phi}\setminus\bigcup_{j=1}^{k}\mathcal{S}_{\phi}^{j}$ is a partition of $\mathcal{S}_{\phi}$. By Proposition 2.1 and Lemma 2.2(4),
$\sum_{j=1}^{k}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}=\sum_{j=1}^{k}\|\mathcal{S}_{\phi}^{j}\|_{s\text{-var}}\leq\|\mathcal{S}_{\phi}\|_{s\text{-var}}<\infty.$
Suppose now that $\sum_{j=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}<\infty$. Let $X_{1},\dots,X_{n}$ be a partition of $\mathcal{S}_{\phi}$, where the arcs $X_{j}$ are enumerated according to the orientation of $\mathcal{S}_{\phi}$, with $0\in X_{n}$. We consider three subsets $P_{1},P_{2},P_{3}$ of the indices set $\{1,\dots,n\}$, and for indices in each $P_{i}$ we prove corresponding estimates.
Estimate 1: Set $P_{1}=\{n\}$, and let $k\in\mathbb{N}$ be the maximal index such that
@@
Then, by definition of $\mathcal{S}_{\phi}^{k}$ and decreasing property of their diameters, by Proposition 2.1 and by Lemma 2.2 we have
@@
$(\operatorname{diam}{X_{j_{1}}})^{s}+\cdots+(\operatorname{diam}{X_{j_{m}}})^{s}$ $\leq\|X_{j_{1}}\|_{s\text{-var}}+\cdots+\|X_{j_{m}}\|_{s\text{-var}}$
$\leq\|X_{j_{1}}\cup\cdots\cup X_{j_{m}}\|_{s\text{-var}}$
@@
Estimate 3: Finally, set $P_{3}=\{1,\dots,n\}\setminus(P_{1}\cup P_{2})$. For each $j\in P_{3}$, there exists minimal $k_{j}\in\mathbb{N}$ such that $X_{j}\cap\mathcal{S}_{\phi}^{k_{j}}\neq\emptyset$. Moreover, if $j,j^{\prime}$ are two distinct such indices, then $k_{j}\neq k_{j^{\prime}}$. For any $j\in P_{3}$, by the decreasing property of $(\operatorname{diam}\mathcal{S}_{\phi}^{m})_{m\in\mathbb{N}}$, we have
@@
$\|\mathcal{S}_{\phi}\|_{s\text{-var}}\leq 3\sum_{n=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{n}))^{s}<\infty,$
and, by Proposition 2.1, $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc. ∎
@@ 4. Hölder exponents between spirals
(4.2) $\alpha\leq\frac{p+q+r+s-pr+qs}{(2+s-r)(1+q)}.$
Theorem 1.1 also provides bounds on the Hölder exponent of maps between such spirals in an implicit way. Recall that there are functions $\phi,\psi:[2\pi,+\infty)\rightarrow(0,\infty)$ that tend to 00 as $t\rightarrow\infty$ with $\mathcal{S}_{p,q}=\mathcal{S}_{\phi}$ and $\mathcal{S}_{r,s}=\mathcal{S}_{\psi}$. Note that it is non-trivial to explicitly determine $\phi$ and $\psi$, due to the implicit relation between arguments of
@@
for some $t\geq 2\pi$, and the modulus $|z_{t}|$. Namely, while the distance of $z_{t}$ from 00 is indeed just $|z_{t}|$, the value $t$ is not always an argument of $z_{t}$, which makes the naive approach of choosing $\phi(t)=|z_{t}|$ generally incorrect for $p\neq q$. However, it is easier to determine $\phi(t_{k})$ and $\psi(t_{k})$ at $t_{k}=k\pi/2$, for integers $k\geq 4$, which is enough to imply that $\mathcal{S}_{\phi},\mathcal{S}_{\psi}$ are almost circular. This can also be seen through the relation of these spirals to the corresponding concentric ellipses (see [BFF22, p. 7]). Moreover, by the aforementioned values $\phi(t_{k})$ and $\psi(t_{k})$, we conclude that
@@
for all $j\in\mathbb{N}$. Suppose that $h:\mathcal{S}_{p,q}\rightarrow\mathcal{S}_{r,s}$ is $\alpha$-Hölder. By Theorem 1.1, there is a $\beta$-Hölder map $g:[0,1]\rightarrow\mathcal{S}_{p,q}$, for all $\beta<p$. Thus, $h\circ g:[0,1]\rightarrow\mathcal{S}_{r,s}$ is $\alpha\beta$-Hölder, which by Theorem 1.1, and the fact that $\beta$ can be as close to $p$ as needed, implies that
@@
Suppose $p>1$ and $r\leq 1$. The latter implies that
$\frac{(2-r)r-1}{1-r}<r\leq s,$
which leads to
$\frac{r(2-r+s)}{1+s}\leq 1\leq p.$
The above inequality is enough to conclude that
$\frac{r}{p}\leq\frac{1+s}{2+s-r},$
which shows that the inequality (4.3) achieved by Theorem 1.1 is indeed an improvement upon the bound (4.1).
The bound (4.3) is also an improvement on the bound stemming from Theorem 4.1 for polynomial spirals with $p=q\in(1,+\infty)$ and $0<r=s\leq 1$, i.e., an improvement on [BFF22, Corollary 2.10]. In fact, this is a sharp bound even for more general positive $p=q$, $r=s$, as stated in Theorem 1.2, which we are ready to prove.
@@
Let $0<r\leq p$. If $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ is $\alpha$-Hölder, then is has already been shown that $\alpha\leq p/r$ in (4.3).
@@
It should be noted that the improved bound (4.3) in the context of Hölder classification for spirals $\mathcal{S}_{p}$, $\mathcal{S}_{r}$ coincides with the sharp bound in the quasiconformal classification problem resolved by Tyson and the first author in [CGT23]. In particular, by [CGT23, Theorem 1.1], two spirals $\mathcal{S}_{p}$ and $\mathcal{S}_{r}$ are $K$-quasiconformally equivalent if, and only if, $K\geq p/r$ (see [V7̈1] for more details on quasiconformal mappings). It is quite interesting that in the case of these spirals, the sharp exponent bound in the Hölder classification programme suggested by Fraser in [Fra21] is essentially attained from the sharp dilatation bounds for the corresponding quasiconformal classification study. This motivates further the question of whether and under what conditions resolving the quasiconformal classification problem for two objects results in the resolution of the corresponding Hölder classification problem. We refer the interested reader to the discussion in [CG25, Section 5] for more details and related results in higher dimensions.

### v3

@@ 1. Introduction
Given a continuous function $\phi:[2\pi,+\infty)\to(0,\infty)$ with $\lim_{t\rightarrow\infty}\phi(t)=0$, we denote by $\mathcal{S}_{\phi}$ the spiral
@@
[…]
@@
Let $0<r\leq p\leq 1$. There is a $r/p$-Hölder map $f:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$, and a Lipschitz map $g:\mathcal{S}_{r}\rightarrow\mathcal{S}_{p}$. Moreover, every map $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ that is $\alpha$-Hölder needs to satisfy $\alpha\leq r/p$.
@@
for all $x_{1},x_{2}\in X$. The number $\alpha$ is called the Hölder exponent of $f$, and the smallest $C>0$ is the Hölder semi-norm of $f$, denoted by $\mathop{\mathrm{H\ddot{o}ld}}\nolimits_{\alpha}f$.
Recall that a metric space $X$ is a metric arc if there is a homeomorphism $f$ mapping the interval [0,1][0,1][ 0 , 1 ] onto the space $X$. Given an interval $I\subset[0,1]$, we say that $f(I)$ is a subarc of $X$. Furthermore, if the interval $I$ has endpoints $a,b\in[0,1]$, we say that $f(I\setminus\{a,b\})$ is the interior of the subarc $f(I)$. If $X$ and $f(I)$ have distinct interiors, we say that $f(I)$ is a proper subarc of $X$. If $f$ is $\alpha$-Hölder for some $\alpha\in(0,1)$, we say that $X$ is an $\alpha$-Hölder arc.
@@
[…]
A classification of Lipschitz arcs (and even more general Lipschitz curves) was given by Ważewski in 1927: an arc is a Lipschitz arc if and only if $\mathcal{H}^{1}(X)<\infty$, and if $X$ is a Lipschitz arc, then there exists a $(2\mathcal{H}^{1}(X))$-Lipschitz parameterization $f:[0,1]\to X$. See [AO17, Theorem 4.4] for the proof. An analogous result for $\frac{1}{s}$-Hölder arcs does not exist as there is no connection between $\frac{1}{s}$-Hölder parameterizations and $\mathcal{H}^{s}$; see for example [BNV19, Proposition 9.8]. In this section we characterize $\frac{1}{s}$-Hölder arcs in terms of the $s$-variation, the analogue of $\mathcal{H}^{1}$ from the Lipschitz arc case.
A partition of a metric arc $(X,d_{X})$ is a finite collection of subarcs $\mathcal{P}=\{X_{1},\dots,X_{n}\}$ with disjoint interiors, and with their union being equal to $X$. For $s\geq 1$, define the $s$-variation of $X$ by
@@
Letting $\epsilon$ go to 0, we obtain the desired inequality.
@@
The continuity of the function $t\mapsto\|f([0,t])\|_{s\textup{-var}}$ follows by [FV10, Proposition 5.8] applied to the given homeomorphism $f$. ∎
@@
[…]
Let $X_{1},\dots,X_{n}$ be a partition of $X$ and for each $i\in\{1,\dots,n\}$ let $I_{i}=f^{-1}(X_{i})$. Then $\{I_{1},\dots,I_{n}\}$ is an interval partition of [0,1][0,1][ 0 , 1 ]. For each $i\in\{1,\dots,n\}$, let $x_{i},y_{i}\in I_{i}$ such that
$d_{X}(f(x_{i})-f(y_{i}))\geq(1-\epsilon)\operatorname{diam}{f(I_{i})}$
@@
$(\operatorname{diam}{f(I_{i})})^{s}\leq(1-\epsilon)^{-s}d_{X}(f(x_{i}),f(y_{i}))^{s}\leq(1-\epsilon)^{-s}H^{s}\operatorname{diam}{J_{i}}.$
@@
Taking supremum over all partitions and letting $\epsilon\to 0$ we obtain $\|X\|_{s\text{-var}}\leq H^{s}$. Since $H>\mathop{\mathrm{H\ddot{o}ld}}\nolimits_{1/s}{f}$ is arbitrary, and $\mathop{\mathrm{H\ddot{o}ld}}\nolimits_{1/s}{f}$ can be chosen as close to $H_{1/s}(X)$ as necessary, this implies that $\|X\|_{s\text{-var}}\leq H_{1/s}(X)^{s}$.
@@
By Lemma 2.2, $\psi$ is an increasing function from [0,1][0,1][ 0 , 1 ] into [0,1][0,1][ 0 , 1 ] and, by Lemma 2.3, $\psi$ is continuous. Since $\psi(0)=0$ and $\psi(1)=1$, it follows that $\psi$ is surjective. By Lemma 2.2(2), it follows that $\psi$ is in fact a homeomorphism. This allows for the definition of $F=f\circ\psi^{-1}:[0,1]\to X$.
[…]
$d_{X}(F(x),F(y))^{s}$ $=d_{X}(f(x^{\prime}),f(y^{\prime}))^{s}$
@@
[…]
$H_{1/s}(X)\leq\mathop{\mathrm{H\ddot{o}ld}}\nolimits_{1/s}{F}\leq\|X\|_{s\text{-var}}^{1/s}<\infty,$
which allows to apply the already proved opposite direction and yields $H_{1/s}(X)=\|X\|_{s\text{-var}}^{1/s}$ as needed. This equality and the definition of $H_{1/s}(X)$ complete the proof. ∎
@@ 3. Hölder rectifiability of spiral arcs
Suppose that $\mathcal{S}_{\phi}$ is almost circular. Fix $j\in\mathbb{N}$ and $s\geq 1$. On the one hand, if $h_{j}:[0,1]\to\mathcal{S}_{\phi}^{j}$ is the constant-speed Lipschitz parametrization map, then for all $x,y\in[0,1]$,
@@
Note that the last inequality above follows by the fact that $\mathcal{S}_{\phi}^{j}$ is closed, the fact that $\phi$ is strictly decreasing, and an application of the triangle inequality. Hence, $H_{1/s}(\mathcal{S}_{\phi}^{j})\simeq\phi_{j}$, and the statement follows directly from Theorem 3.1. ∎
@@
Fix $s\geq 1$. We show that
(3.1) $\frac{1}{2}\|\mathcal{S}_{\phi}\|_{s\text{-var}}\leq\sum_{j=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}\leq\|\mathcal{S}_{\phi}\|_{s\text{-var}}$
and Theorem 3.1 follows directly from Proposition 2.1.
For the upper bound of (3.1) fix $k\in\mathbb{N}$ and note that $\mathcal{S}_{\phi}^{1},\dots,\mathcal{S}_{\phi}^{k},\mathcal{S}_{\phi}\setminus\bigcup_{j=1}^{k}\mathcal{S}_{\phi}^{j}$ is a partition of $\mathcal{S}_{\phi}$. By Proposition 2.1 and Lemma 2.2(4),
$\sum_{j=1}^{k}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}=\sum_{j=1}^{k}\|\mathcal{S}_{\phi}^{j}\|_{s\text{-var}}\leq\|\mathcal{S}_{\phi}\|_{s\text{-var}}.$
For the lower bound of (3.1), let $X_{1},\dots,X_{n}$ be a partition of $\mathcal{S}_{\phi}$, where the arcs $X_{j}$ are enumerated according to the orientation of $\mathcal{S}_{\phi}$, with $0\in X_{n}$. Let $k\in\mathbb{N}$ be the maximal index so that
@@
We consider three subsets $P_{1},P_{2},P_{3}$ of the indices set $\{1,\dots,n\}$, and for indices in each $P_{i}$ we prove corresponding estimates.
Estimate 1: Set $P_{1}=\{X_{n}\}$. By definition of $\mathcal{S}_{\phi}^{k}$ and decreasing property of their diameters, by Proposition 2.1 and by Lemma 2.2 we have
@@
$(\operatorname{diam}{X_{j_{1}^{l}}})^{s}+\cdots+(\operatorname{diam}{X_{j_{m}^{l}}})^{s}$ $\leq\|X_{j_{1}^{l}}\|_{s\text{-var}}+\cdots+\|X_{j_{m}^{l}}\|_{s\text{-var}}$
$\leq\|X_{j_{1}^{l}}\cup\cdots\cup X_{j_{m}^{l}}\|_{s\text{-var}}$
@@
Estimate 3: Finally, set $P_{3}=\{1,\dots,n\}\setminus(P_{1}\cup P_{2})$. For each $j\in P_{3}$, there exists a minimal $k_{j}\in\mathbb{N}$ such that $X_{j}\cap\mathcal{S}_{\phi}^{k_{j}}\neq\emptyset$. Moreover, if $j,j^{\prime}$ are two distinct such indices, then $k_{j}\neq k_{j^{\prime}}$. For any $j\in P_{3}$, by the decreasing property of $(\operatorname{diam}\mathcal{S}_{\phi}^{m})_{m\in\mathbb{N}}$, we have
@@
$\|\mathcal{S}_{\phi}\|_{s\text{-var}}\leq 2\sum_{n=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{n}))^{s}.\qed$
@@ 4. Hölder exponents between spirals
(4.2) $\alpha\leq\frac{p+q+r+s-pr+qs}{(2+s-r)(p+q)}.$
For $p\leq 1$, Theorem 1.1 also provides bounds on the Hölder exponent of maps between such spirals in an implicit way. Recall that there are functions $\phi,\psi:[2\pi,+\infty)\rightarrow(0,\infty)$ that tend to 0 as $t\rightarrow\infty$ with $\mathcal{S}_{p,q}=\mathcal{S}_{\phi}$ and $\mathcal{S}_{r,s}=\mathcal{S}_{\psi}$. Note that it is non-trivial to explicitly determine $\phi$ and $\psi$, due to the implicit relation between arguments of
@@
for some $t\geq 2\pi$, and the modulus $|z_{t}|$. Namely, while the distance of $z_{t}$ from 0 is indeed just $|z_{t}|$, the value $t$ is not always an argument of $z_{t}$, which makes the naive approach of choosing $\phi(t)=|z_{t}|$ generally incorrect for $p\neq q$. However, it is easier to determine $\phi(t_{k})$ and $\psi(t_{k})$ at $t_{k}=k\pi/2$, for integers $k\geq 4$, which is enough to imply that $\mathcal{S}_{\phi},\mathcal{S}_{\psi}$ are almost circular. This can also be seen through the relation of these spirals to the corresponding concentric ellipses (see [BFF22, p. 7]). Moreover, by the aforementioned values $\phi(t_{k})$ and $\psi(t_{k})$, we conclude that
@@
for all $j\in\mathbb{N}$. Suppose that $h:\mathcal{S}_{p,q}\rightarrow\mathcal{S}_{r,s}$ is $\alpha$-Hölder. By Theorem 1.1, there is a $\beta$-Hölder map $g:[0,1]\rightarrow\mathcal{S}_{p,q}$, for all $\beta<p$. Thus, the map $h\circ g:[0,1]\rightarrow\mathcal{S}_{r,s}$ is $\alpha\beta$-Hölder, which by Theorem 1.1, and the fact that $\beta$ can be as close to $p$ as needed, implies that
@@
For polynomial spirals studied in [BFF22], the above bound is always an improvement upon Theorem 4.1. In particular, for $p=q$ and $r=s$ the bound (4.2) implies that
$\alpha\leq\frac{p+r}{2p}.$
In the case where $p\geq r$, the exponent bound (4.3) is indeed an improvement on the above bound stemming from Theorem 4.1, i.e., an improvement on [BFF22, Corollary 2.12]. In fact, this is a sharp bound in a more general sense, as stated in Theorem 1.2, which we are ready to prove.
@@
Let $0<r\leq p$. If $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ is $\alpha$-Hölder, then is has already been shown that $\alpha\leq r/p$ in (4.3).
@@
Despite the restriction to polynomial spirals for sharpness, our bound (4.3) can be shown to be an improvement upon Theorem 4.1 even for elliptical spirals with specific values of $p\neq q,r\neq s$. In particular, if $0<r<p\leq 1$, then we have the strict inequality
$\frac{r}{p}<\frac{p+r}{2p}.$
This gap between the two values, along with the continuity of the expression
$\frac{p+q+r+s-pr+qs}{(2+s-r)(p+q)}$
in $q$ and $s$, implies that for $q,s$ close enough to $p,r$, respectively, our bound (4.3) is an improvement upon (4.2) for elliptical, non-polynomial spirals as well. For instance, if $p=3/5$, $r=s=1/2$, we have that the exponent bound (4.3) is an improvement upon (4.2) for certain values of $q$ close to 3/53/53 / 5, i.e.,
$\frac{r}{p}=\frac{5}{6}<\frac{13+15q}{12+20q}=\frac{p+q+r+s-pr+qs}{(2+s-r)(p+q)},$
for all $q\in(3/5,9/5)$. However, it is clear that this is not generally the case, as can be seen by choosing any $q>9/5$ in the above inequality.
It should also be noted that the improved bound (4.3) in the context of Hölder classification for spirals $\mathcal{S}_{p}$, $\mathcal{S}_{r}$ coincides with the sharp bound in the quasiconformal classification problem resolved by Tyson and the first author in [CGT23]. In particular, by [CGT23, Theorem 1.1], two spirals $\mathcal{S}_{p}$ and $\mathcal{S}_{r}$ are $K$-quasiconformally equivalent if, and only if, $K\geq p/r$ (see [V7̈1] for more details on quasiconformal mappings). It is quite interesting that in the case of these spirals, the sharp exponent bound in the Hölder classification programme suggested by Fraser in [Fra21] is essentially attained from the sharp dilatation bounds for the corresponding quasiconformal classification study. This motivates further the question of whether and under what conditions resolving the quasiconformal classification problem for two objects results in the resolution of the corresponding Hölder classification problem. We refer the interested reader to the discussion in [CG25, Section 5] for more details and related results in higher dimensions.